Please help me solves these urgent problems!!!
first system
1. x^2+y^2=16
2x+y=4


second system
2. x^2+y^2=4
y=x^2-2

Answer by Alan3354(69443)

first system
1. x^2+y^2=16
2x+y=4 --> y = -2x+4
Sub for y into the 1st eqn
x^2 + (-2x+4)^2 = 16
x^2 + 4x^2 -16x + 16 = 16
5x^2 - 16x = 0
x = 0
x = 16/5
---------
Solve for y on the line, not the circle, to avoid extraneous roots
x = 0, y = 4 --> (0,4)
x = 16/5, y = -12/5 --> (3.2,-2.4)
-----------------------
second system
2. x^2+y^2=4
y=x^2-2 --> x^2 = y+2
Sub for x^2
(y+2) + y^2 = 4
y^2 + y - 2 = 0
(y+2)*(y-1) = o
y = 1, y = -2
---------------
x^2 = y+2
x^2 = 0 --> (0,-2)
-------------------
x^2 = 3 --> (sqrt(3),1) and (-sqrt(3),1)